Can someone help me out with this formula with an example?

VAR(?? ) = 1 ? − 1 ∑[??,? − ?(?? )] 2 ? ?=1 ,

Thank you
Answer #1

Lets say

there is 3 year data
T=3

Year Returns
1 10%
2 15%
3 20%

E(Ri)=(10%+15%+20%)/3=15%

VAR(Ri)=1/(3-1)*((10%-15%)^2+(15%-15%)^2+(20%-15%)^2)=0.00250